Displaying reports 1061-1079 of 1079.Go to page Start 50 51 52 53 End
Reports until 19:26, Thursday 15 June 2023
Controls Optimal Controls
Ian MacMillan - posted 19:26, Thursday 15 June 2023 - last comment - 15:09, Friday 16 June 2023(11198)
Cleaned the Buzz Repo

I have cleaned up the buzz repo and made sure to update the solvers.py file with the new BH1_solvers_test.py from the old 300m repo. I have rerun the ASC notebooks and verified that the new code still functions properly. I have tagged this version as "cleaned_v1".

I also added pytests in the file T_Test4Lee. These tests will allow for tests to be run on the currently working solver as well as developing a new solver in  the BH1_solvers_test.py file. By setting the solver in the calcOpt function as "test" the solver will use the solver in the file. This can also be done by setting the solver in the test params to "test". Hopefully this allows for easy prototyping of the solver with out interfering with the other solvers or code

Comments related to this report
Ian MacMillan - 15:09, Friday 16 June 2023 (11201)

Gitlab CI now works. I had to disable the LSC and DAMP tests to get it to run but the ASC one is the only one that I am looking at currently.

Controls Optimal Controls
Ian MacMillan - posted 15:48, Wednesday 14 June 2023 (11197)
Buzz code matching old 300m cavity code

[Ian, Lee]

Using saved versions of the models from the 300m cavity repo (https://git.ligo.org/wieldphysics/ligo300mfiltercavity) we have matched the resulting controllers from the 'ligo300mfiltercavity/T_ASC_mixed_H_BH2.py solver2' test with new buzz code (https://git.ligo.org/ian.macmillan/buzz).

The 300m result shows the mixed H2/Hinf result at one weight for the two FOMs: (BNS FOM: 250000, and the Flat FOM: 1). The multiple traces comes from the gamma parameter being decreased flattening the closed loop gain. These results (with the file prefix "300m") are in the attachments. These controllers match with the open and closed loop gain that we get from the Buzz code also in the attachments with the prefix "buzz".

I also attached two RMS Plots. The First one (with the prefix "RMS_single" is of the single FOM weighting which is the controllers we calculated in the comparison above. As is seen on the plot the phase margin diverges from zero as the gamma decreases. This is what we expect but was not what we were seeing earlier because of a bug. It may be wise not to trust 'control.margin()'.

Once we know that the result of our two codes match I wanted to explore the what other weightings did. The last plot (prefix "RMS_multi") is using a weighting of F1_gain form 250 to 250000 with 4 different weightings, where F2_gain (flat FOM) is 1/F1_gain. The results show that either the solver is very fragile or maybe I haven't picked the correct weighting. It is possible that the weighting is well out of the stable regime for the solver. This would be strange because the max F1 weighting is 250,000 which should be the most unstable weighting combo since the solver needs a non-zero D matrix. This is disappointing but one good thing is that its border seems to follow the nice optimal RMS limiting curve set by the H2 solver.

Non-image files attached to this report
Lasers General
Torrey Cullen - posted 11:43, Wednesday 14 June 2023 - last comment - 14:42, Thursday 22 June 2023(11196)
First acquired lock on the cryo lab west cavity

I successfully locked the cryo west cavity using the laser lock box on the moku yesterday. The main catalyst to this was that the PD being used to measure the REFL signal had a bandwidth of ~10 MHz and was attenuating the 32.7 MHz modulation signal. We ordered some newport 1811 PDs (BW = 125 MHz), and the modulation signal shot up by several orders of magnitude. Note that there is now a lens in front of the REFL PD because the diode on the 1811 is very small, also the damage threshold on these is MUCH LOWER than some of the thorlab equivalents. Don't send > 1mW towards the cavity.

First, the settings of the laser lock box that were used for initial lock can be seen in the attached text file, as well as a screen shot of the lockbox while locked. The process I was using was:

1) Make sure the TEC setpoint is roughly centered to ensure maximum range on the slow loop. Manual engage the slow loop.

2) Slide the slow loop offset until its roughly near the cavity resonance, i.e. you see the large dip in REFL PD.

3) Engage the laser lock box assist function, which is this button. You should now see the REFL PD dip is centered on the error signal zero crossing point, along with some cross-hairs highlighting the zero crossing point on the error signal. These are clickable. Click the zero point that corresponds to the RELF PD dip (sometimes the moku adds additional cross-hairs that are garbage). I'm realizing I don't have a screenshot of this, I'll grab one and comment on this post with it later today.

And thats it! Pretty easy with the moku once everything is set up. I kept the cavity locked for 10+ minutes while tapping the table, the lock seems fairly robust. A minor nuisance, in the other attached photo there seems to be some random low frequency oscillations on the REFL PD that I'm unsure where its coming from. 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 13:23, Friday 16 June 2023 (11200)

Example of zero point crossing when using the lock assist function.

Images attached to this comment
Ian MacMillan - 15:14, Friday 16 June 2023 (11202)

Laser's steady hum,

Cryo cavity now locked,

Early victory.

Torrey Cullen - 14:42, Thursday 22 June 2023 (11209)

Diagram of current setup when lock was achieved. 

Images attached to this comment
Non-image files attached to this comment
GQuEST Meeting
Chris Stoughton - posted 10:25, Tuesday 13 June 2023 (11195)
Initial test of meeting notes
This is a log entry for meetings. Let's use Section=GQuEST and Primay Task=Meeting for these.
Controls General
Ian MacMillan - posted 17:20, Tuesday 06 June 2023 (11193)
LSC Model Not Playing Nice with Mixed H2/Hinf solver

I tried the LSC model (tried attaching but does not accept .mat files) with the H2 solver And it worked like normally. You can see these results in the "LSC_H2_Results.pdf" attachment.

When trying the LSC model with the H2/Hinf Mixed solver it struggled to recover even the H2 results. This is with a gamma going from 1e3 down to 1e-1 (not igsq). On the H2 result it looks like the Closed loop is peaking at about 100 so this should be plenty to at least recover the H2 results. but I am left with the results seen in "LSC_H2HinfMix_Results.pdf"

It is possible that this is a problem with the Zinf connection but I will have to investigate further. I also still need to try the damping loop. Also I got multiple color bars working using This.

Non-image files attached to this report
Controls General
Ian MacMillan - posted 13:38, Monday 05 June 2023 - last comment - 12:49, Tuesday 06 June 2023(11190)
Paper Plots for H2/Hinf Optimal Controls Paper
Here I make a list of all of the plots that I want to include in the final $H_2$/$H_{inf}$ Optimal Controls paper. The paper should look at 3 models: - The damping loop - The LSC - The ASC For each of these models should have a few plots: 1) Open Loop Plot: This plot should show the open loop transfer function of the full system. It should show the current loop as well as a number of different loops for varying FOM weight and varying $H_{inf}$ upper limit {gamma} 2) An RMS Plot: This will show the $H_2$ norm (RMS output) from the white noise inputs to the FOM outputs. This plot will be a scatter plot with one FOM being the BNS FOM and one being the flat FOM 3) Closed Loop Plot: Although this may not make it into the paper I need to make a closed loop plot that shows the closed loop transfer function for the current controller and the controllers calculated for decreasing $H_{inf}$ upper limit {gamma}. This plot will show the mixed controller is limiting the peak of the closed loop magnitude. 4) A Noise Plot: this plot will characterize the plant model for the given models above it will include the magnitude of the noise blocks. In addition to all those plots I will want to create some general plots that are one off: 1) I will want to create a FOM plot which shows the Flat FOM the BNS FOM and the simplified BNS FOM that is used. Only Magnitude will be necessary
Comments related to this report
Ian MacMillan - 12:49, Tuesday 06 June 2023 (11192)

I am running into a problem choosing the color scheme for the plots for the paper. There are two main challenges to this first is that in the mixed H2/Hinf case the color conveys two pieces of info: the FOM weighting and the gamma limit on the Hinf norm. I have attached a number of single color schemes in the "PaperALS_OL_BODE_colors_comp.pdf" attachment. I don't know how to handle the gamma limit. Ideally there would be a 2d color space to map like in the pycolormap-2d package. Is it worth it?

The other main plot is the RMS plot which also has two color schemes one for the gain margin and one for the FOM weighting. This is easier as long as the colors don't overlap. Also I haven't gotten the dual color bar to work on the RMS Plot. the scaling always is removed by adding the second color bar. See the example RMS plot in "PaperALS_FOM_RMS_Scatter.pdf"

Note: some of the colorbars are messed up in "PaperALS_OL_BODE_colors_comp.pdf"

Non-image files attached to this comment
Lasers General
Torrey Cullen - posted 10:42, Friday 02 June 2023 - last comment - 08:34, Monday 05 June 2023(11188)
Frequency noise measurement

Remeasured the frequency noise with the 785 after the visibility improvement. Assuming my calibration to Hz/sqrt(hz) is correct, the measurement is still yielding several orders of magnitude larger than what rio claims it to be. Investigating this today.

 

For reference, the SR785 measures a quantity is V/sqrt(Hz). To convert to frequency noise I do S / (L/c * B*2*pi) where S is the measurement, L is the path length difference between the two arms and B is the fringe amplitude. V/sqrt(Hz) / (m/(m/s) * V) = Hz/sqrt(Hz).

Images attached to this report
Comments related to this report
Lee McCuller - 08:34, Monday 05 June 2023 (11189)

The right axis (Hz) on this plot has a different frequency dependence than the left axis (Rad). So be careful with the specification data. I agree otherwise that there seems to be an excess. It might be intensity noise that needs to be canceled.

Lasers General
Torrey Cullen - posted 09:35, Friday 02 June 2023 - last comment - 09:35, Friday 02 June 2023(11185)
Laser Frequency Characterization with Mokus

Lee came into the lab today and discovered that the TEC controller's range on the integrator is very small, which is why I was having trouble driving the laser frequency. If you center on this range (~10.2 kOhm) you can then drive a signal into the back of the TEC so that there are constant fringes on the Mach-Zehnder read out. I did this via the moku (15 mV @ .1 Hz, corresponds to roughly 1 fringe). "2.png" is this first attempt, red is the readout PD, blue is the driving signal. "3.png" is then hooking up a PD farther upstream to use as a power reference instead of the driving signal (see also cryolab_frequency_noise_measurement_pdf.pdf, the lighter color laser path in that is used). "4.png" is then using the moku's math channels to do (readout PD)/(power reference) to give the fringes at a constant offset. At this point its very easy to align to maximize these fringes. "7.png" shows the improved visibility. "5.png" shows the moku set up used. There are some shenanigans with the input ranges when using an oscilloscope in multi-instrument mode where you either have it on 400mV setting and its saturated, or its on 4V setting and it auto attenuates the signal by 20dB. So the digital filter box is undoing that 20dB attenuation. For unequal powers, I believe the formula is V =(2*sqrt(34*99)/(99+34))=87%, so not perfect but a vast improvement.

Images attached to this report
Non-image files attached to this report
Comments related to this report
Torrey Cullen - 17:00, Thursday 01 June 2023 (11186)
Also, how do you link to images on this. I know you can on the LLO logbook. So I don't have to keep writing "*.png".
Lee McCuller - 09:34, Friday 02 June 2023 (11187)

I drag and drop the link at the bottom of the post during editing and it should put it inline in the text. - looks like you have to be in the ckeditor for that to work well.

macmillan@caltech.edu - posted 15:32, Tuesday 30 May 2023 - last comment - 11:21, Thursday 01 June 2023(11176)
Hello World!
According to all known laws of aviation, there is no way a bee should be able to fly. Its wings are too small to get its fat little body off the ground. The bee, of course, flies anyway because bees don't care what humans think is impossible. Yellow, black. Yellow, black. Yellow, black. Yellow, black. Ooh, black and yellow! Let's shake it up a little. Barry! Breakfast is ready! Coming! Hang on a second. Hello? Barry? Adam? Can you believe this is happening? I can't. I'll pick you up. Looking sharp. Use the stairs, Your father paid good money for those. Sorry. I'm excited. Here's the graduate. We're very proud of you, son. A perfect report card, all B's. Very proud. Ma! I got a thing going here. You got lint on your fuzz. Ow! That's me! Wave to us! We'll be in row 118,000. Bye! Barry, I told you, stop flying in the house! Hey, Adam. Hey, Barry. Is that fuzz gel? A little. Special day, graduation. Never thought I'd make it. Three days grade school, three days high school. Those were awkward. Three days college. I'm glad I took a day and hitchhiked around The Hive. You did come back different. Hi, Barry. Artie, growing a mustache? Looks good. Hear about Frankie? Yeah. link [test](google.com)
Comments related to this report
macmillan@caltech.edu - 16:09, Tuesday 30 May 2023 (11177)
i disagree
macmillan@caltech.edu - 16:10, Tuesday 30 May 2023 (11178)
i agree too
Alex Ramirez - 11:21, Thursday 01 June 2023 (11182)
To agree is to disagree to disagreeing
admin - posted 15:18, Friday 21 May 2010 (1)
The OSLogbook
Welcome to the OSLogbook application.
Displaying reports 1061-1079 of 1079.Go to page Start 50 51 52 53 End